OpenStudy (anonymous):

What is 5263 × 10–3 ? A. 5.263 B. 52.63 C. 526.3 D. 526,300

OpenStudy (anonymous):

\[5263x10^{-3} =\frac{ 5263 }{ 10^3 } = \frac{ 5263 }{ 10*10*10 }\]

OpenStudy (anonymous):

can you solve now ^^

OpenStudy (anonymous):

Yes thank you

OpenStudy (phi):

move the decimal point 3 places to the left.

OpenStudy (anonymous):

52627